# Predicting perturbation targets with causal differential networks

### v1

@@ 3.1 Causal featurizer
    Given a dataset $D$, smaller batches are constructed by sub-sampling both examples and variables. Heuristics like pairwise correlation are used to select variables which are likely to have causal relationships, to minimize unnecessary computation.
@@
    Two sets of input features are generated from these batches: global summary statistics like correlation, computed over all variables, and marginal estimates, the outputs of classical causal discovery algorithms (Spirtes et al., 1995) run on small subsets of variables.
@@
    Finally, these two sets of features are input to a neural network (“aggregator”), which was trained to map them into causal graphs on synthetic datasets.
We run this procedure on $D_{\text{obs}}$ and $D_{\text{int}}$ independently, treating both as “observational” datasets. To obtain the graph representations for each dataset, we extract the aggregator’s last layer hidden representations before they are collapsed into binary edge predictions, yielding $h_{\text{obs}}$ and $h_{\text{int}}$. Finally, we concatenate $h_{\text{obs}}$ and $h_{\text{int}}$ along the hidden dimension, to obtain a paired graph representation of size $N\times N\times 2d$.
Figure 2: Model overview. The observational and interventional datasets are featurized using a pretrained, amortized causal discovery model. The pair of graph representations are concatenated and input to an axial attention model. After the 2D attention, representations are aggregated over incoming edges to obtain binary, node-level predictions. Cdn is trained with a classification objective, over synthetic and/or real data. Gray modules are frozen. Yellow modules are learned.
3.2 Differential network
Given the graph representation, the differential network predicts which nodes were intervened upon (Figure 2, right). Its architecture is composed of an axial-attention layer, followed by a linear projection. Following Sea, the axial attention layer attends separately along the rows and columns of the graph’s adjacency matrix. This operation is equivalent to self-attention along all nodes in the “outgoing edge” direction, with the “incoming edges” as a batch dimension, followed by the opposite. We use pre-layer normalization on each self-attention, followed by dropout and residual connections:
$h=h+\text{Dropout}(\text{Self-Attn}(\text{LayerNorm}(h))).$ (3)
After the axial attention, we mean over all “incoming edges” and linearly project to binary node-level predictions, where a normalized 1 indicates an intervention target, and 0 indicates otherwise.
3.3 Implementation and training details
For the causal featurizer, we used a frozen, pretrained aggregator from (Wu et al., 2024), which corresponded to the FCI-based marginal estimates (Spirtes et al., 1995). The differential network was implemented with one axial attention layer, following the same architecture as the pretrained aggregator, with twice the model dimension (paired graphs). We swept over the number of layers (1-4) on synthetic data and found that a single layer performed the best. In terms of model design, we ablated replacing the attention architecture with a multi-layer perceptron (no communication between nodes, Table 4), as well as other components (Table 7) and summary statistics (Table 8). Due to the large graph sizes (adjacency matrices up to $1000^{2}$), standard 1D attention was intractable. Following Sea, we used inverse covariance as the global statistic on synthetic data, and correlation on transcriptomics data.
We trained separate differential networks on synthetic and real data. On synthetic data, we used a batch size of 16 and the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 1e-4 (identical to Sea). On real data, we started from the synthetic Cdn checkpoint, and then used a batch size of 2 with a learning rate of 1e-5, due to the larger graph sizes. All models took around 4-10 hours to train on a single A6000 GPU.
3.4 Theoretical context
This paper primarily describes a causality-inspired approach towards perturbation target prediction. Contrary to causal discovery algorithms that are self-contained and provably sound under standard assumptions, we cannot “prove” that a pretrained model extracts correct graphs on real data, as the true graphs are unknown. Our evaluation focuses solely on perturbation target prediction, as the ground truth is known, by (experimental) design. To contextualize the predictions of our model, we provide brief sketches with regards to our model’s capacity to map paired graph features to intervention targets. We defer formal claims to Appendix A.
Claim 3.1 (Hard interventions, informal).
Given a causal graphical model $G=(V,E)$ and intervention targets $I\subseteq V$, let $E^{\prime}$ denote the adjacency matrix of mutilated graph $G^{\prime}$ after hard intervention on $I$. An axial attention layer is well-specified to map $(E,E^{\prime})$ to $I$.
Claim 3.2 (Soft interventions, informal).
Let $G$ be a graphical causal model associated with data distribution $P_{X}$, and let $\tilde{P}_{X}$ be the interventional distribution after soft intervention on $I\subsetneq V$. Let $R,R^{\prime}$ denote the correlation matrices and let $\Sigma,\Sigma^{\prime}$ denote the covariance matrices of $X,\tilde{X}$. An axial attention layer is well-specified to map $(R,R^{\prime},\Sigma,\Sigma^{\prime})$ to $I$.
4 Experiments and results
Table 1: Final biological dataset statistics. “gw” and “es” are two independent screens with different perturbations.
Type Cell line ##\## Perts ##\## Cells
Genetic K562 gw 1767 492,096
K562 es 1060 213,552
RPE1 961 179,696
HepG2 627 79,309
Jurkat 1014 174,698
Chemical A172 3 18,196
T98G 3 13,126
Figure 3: Data illustration and splitting procedure.
@@ 4 Experiments and results
We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1). Note that these baselines all use various domain knowledge to inform their predictions. For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
@@ 4.1 Biological experiments
We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al. (2022) and Nadig et al. (2024); as well as two Sci-Plex (Srivatsan et al., 2020) datasets (chemical perturbations) from McFaline-Figueroa et al. (2024). Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$ (Figure 3A). In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target.
To ensure high quality labels, we filtered perturbations to those that induced over 10 differentially-expressed genes (statistically significant change, compared to control), of which the true target should be present. This is to exclude perturbations with insufficient cells (low statistical power), with minimal to no effect (uninteresting), and those that did not achieve the desired effect (CRISPR efficacy is not guaranteed). For evaluation, we limited the set of candidate targets to the top 1000 differentially expressed genes, ranked by log-fold change per perturbation. If fewer genes were differentially expressed, we randomly sampled additional candidates until we reached a minimum of 100 genes. Finally, we also stratify genetic perturbations based on whether the target is trivially identifiable as the gene with the largest log-fold change (“trivial” vs. “non-trivial”). This is because genetic perturbations are highly specific by design, but also constitute the largest single-cell perturbation datasets. The final data statistics are shown in Table 1, with additional details in Appendix B.1.
Evaluation
@@
We consider two splits: seen and unseen cell lines (Figure 3B). In the former, models may be trained on approximately half of the perturbations from each cell line, and are evaluated on the unseen perturbations. In the latter, we hold out one cell line at a time, and models may be trained on data from the remaining cell lines. Note that not all baselines can be evaluated on unseen cell lines. To ensure that our train and test splits are sufficiently distinct, we cluster perturbations based on their log-fold change and assign each cluster to the same split.
@@
While perturbation target prediction appears to be a simple classification task, there are two aspects of the data that render standard metrics less meaningful. Not all genes can be mapped to the baselines’ domain knowledge, so their effects as perturbations cannot be predicted. In addition, due to genetic redundancy, it is common for multiple perturbations to elicit similar responses (Kernfeld et al., 2023). Therefore, an “incorrectly” predicted target (e.g. out of 1000 genes) may not necessarily indicate poor performance. Therefore, we propose the following metrics.
@@
  • We report the rank of the ground truth target, normalized by the number of candidate genes. Here, 1 indicates that the ground truth target is at the top of the list (best), while 0 indicates that the ground truth is at the bottom (worst).
@@
  • To emulate a “virtual screening” setting, we report recall at k (the proportion of targets recovered within the top $k$ candidates). This value ranges from 0 to 1, where 1 is best.
  • The Pearson correlation $r$ measures the similarity between the (ground truth) mean log-fold change of the top prediction that was observed as a perturbation, and that of the actual perturbation (given as input). These values range from -1 to 1, where 1 is best.
Baselines
We compare to two classes of algorithms related to biological perturbations. Forward inference models predict the effects of perturbations, and perturbation targets can be ranked by comparing the inferred effect of each candidate to the interventional data.
  • Gears (Roohani et al., 2023) is a graph neural network that predicts the effects of unseen genetic perturbations, using the gene ontology knowledge graph (Ashburner et al., 2000) to model the relationship between genes. We trained Gears on each cell line separately, and predicted the effects of perturbing every expressed gene. Then we ranked candidates based on cosine similarity to the interventional data.
  • GenePt (Chen & Zou, 2023) is a set of natural language-based embeddings that have been shown to achieve state-of-the-art performance on unseen perturbation effect prediction with simple downstream models (Märtens et al., 2024). We concatenated each perturbation’s gene embedding to the log-fold change of the top 5000 highly-variable genes (Wolf et al., 2018) and trained a logistic regression model on each cell line to predict true vs. decoy perturbations. Candidates were ranked based on predicted probability.
Predictive models directly predict perturbation targets from data as a classification task.
  • Linear and Mlp take as input the mean expression of all perturbation targets, plus the top 2000 highly-variable genes (Wolf et al., 2018). They are trained to predict a binary label for each gene on each cell line independently.
  • PDGrapher (Gonzalez et al., 2024) is a graph neural network that predicts the perturbation targets of genetic or chemical perturbations, on seen or unseen cell lines, with the human reference interactome (Luck et al., 2020) as the knowledge graph.
@@
All baselines were run with their official implementations and/or latest releases. For more details, please see Appendix C.1.
@@
Table 2: Results on 5 Perturb-seq datasets. Top: “trivial” perturbations. Bottom: “non-trivial” perturbations. Primary setting: train all cell lines jointly, test on unseen perturbations. Suffix -cl: leave out the one cell line from training, test on the same set of unseen perturbations in the unseen cell line. Metrics, from left to right: normalized rank of ground truth, top 20 recall, and top 1 Pearson correlation. For ablations, runtimes, and uncertainties, see Appendix D.
K562 gw K562 es RPE1 HepG2 Jurkat
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Random $0.50$ $0.13$ $0.16$ $0.50$ $0.09$ $0.17$ $0.50$ $0.09$ $0.36$ $0.50$ $0.06$ $0.34$ $0.50$ $0.08$ $0.22$
Linear $0.49$ $0.16$ $0.21$ $0.50$ $0.13$ $0.31$ $0.48$ $0.09$ $0.45$ $0.48$ $0.07$ $0.43$ $0.48$ $0.08$ $0.29$
Mlp $0.48$ $0.21$ $0.20$ $0.47$ $0.13$ $0.21$ $0.54$ $0.19$ $0.37$ $0.51$ $0.10$ $0.42$ $0.48$ $0.13$ $0.32$
GenePt $0.60$ $0.39$ $0.39$ $0.47$ $0.11$ $0.26$ $0.57$ $0.33$ $0.55$ $0.44$ $0.12$ $0.41$ $0.55$ 0.32 0.53
Gears $0.49$ $0.14$ $0.18$ $0.51$ $0.09$ $0.27$ $0.49$ $0.09$ $0.40$ $0.48$ $0.09$ $0.40$ $0.47$ $0.06$ $0.24$
PdG $0.49$ $0.20$ $0.26$ $0.52$ $0.18$ $0.34$ $0.51$ $0.11$ $0.42$ $0.46$ $0.06$ $0.40$ $0.49$ $0.11$ $0.35$
Cdn 0.77 0.52 0.49 0.68 0.32 0.45 0.74 0.37 0.55 0.68 0.20 0.50 0.67 $0.27$ $0.45$
PdG-Cl $0.40$ $0.04$ $0.15$ $0.41$ $0.02$ $0.24$ $0.48$ $0.02$ $0.37$ $0.47$ $0.03$ $0.41$ $0.53$ $0.06$ $0.33$
Cdn-Cl 0.71 0.43 0.40 0.66 0.27 0.42 0.62 0.24 0.53 0.66 0.17 0.46 0.65 0.26 0.47
Random $0.50$ $0.11$ $0.17$ $0.50$ $0.04$ $0.17$ $0.50$ $0.04$ $0.44$ $0.50$ $0.04$ $0.33$ $0.50$ $0.06$ $0.21$
Linear $0.51$ $0.18$ $0.23$ $0.48$ $0.01$ $0.25$ $0.49$ $0.03$ $0.52$ $0.51$ $0.04$ 0.43 $0.50$ $0.06$ $0.23$
Mlp $0.45$ $0.19$ $0.21$ $0.49$ $0.04$ $0.11$ $0.53$ $0.09$ $0.41$ $0.48$ 0.07 $0.40$ $0.45$ $0.07$ $0.29$
GenePt $0.52$ $0.24$ $0.32$ $0.42$ $0.03$ $0.19$ $0.41$ $0.06$ $0.46$ $0.31$ $0.03$ $0.40$ $0.51$ 0.15 $0.36$
Gears $0.56$ $0.16$ $0.22$ $0.45$ $0.06$ $0.20$ $0.49$ $0.04$ $0.44$ $0.54$ $0.03$ $0.40$ $0.50$ $0.11$ $0.26$
PdG $0.54$ $0.24$ $0.29$ $0.49$ $0.03$ $0.23$ $0.48$ $0.05$ $0.52$ $0.54$ $0.05$ 0.43 $0.52$ $0.10$ $0.35$
Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 $0.40$ 0.62 $0.10$ 0.39
PdG-Cl $0.39$ $0.02$ $0.13$ $0.44$ $0.01$ 0.28 0.54 $0.02$ 0.48 $0.48$ $0.02$ 0.41 $0.57$ $0.06$ $0.35$
Cdn-Cl 0.68 0.25 0.29 0.55 0.06 $0.27$ $0.49$ 0.07 0.48 0.59 0.05 $0.39$ 0.60 0.18 0.38
Figure 4: Recall at $k$ on K562 datasets.
@@
Table 3: Rank of drug targets on chemical perturbation datasets (McFaline-Figueroa et al., 2024). A172 and T98G are unseen cell lines. For details and full drug names, please see Appendix B.1.
@@
PdG $0.719$ $0.346$ $0.395$ $0.055$ $0.345$ $0.420$
Cdn 0.883 0.436 0.582 0.059 0.543 0.647
Results
Table 2 reports our results on five Perturb-seq datasets. In the majority of cases, Cdn outperforms baselines, both in the seen and unseen cell line settings. This is more evident in Figure 4, in which Cdn achieves higher recall at $k$ at nearly all points on the curves. While no baseline is as consistent as Cdn, GenePt is surprisingly competitive against more complex baselines. This may indicate that natural language induces embedding spaces with favorable geometries with respect to biological function. On the two chemical perturbation datasets, Cdn ranks the ground truth targets higher than PdGrapher on all six drugs (Table 3). However, the mutually low performance on doxorubicin (doxo.) in T98G cells may indicate a failure mode of both models.

### v2

@@ 3.3 Implementation details
    Initialize the causal structure learner with pretrained weights from Wu et al. (2024) (except ablations).
@@
    Train both modules jointly on synthetic data.
@@
    Finetune both modules jointly with only $\mathcal{L}_{I}$ (no ground truth graphs), on each training split of the real datasets.
For the synthetic training data, we generated 8640 datasets with hard and soft interventions, which vary in causal mechanism, intervention type, and graph topology. To emulate transcriptomics data, in which perturbation effect sizes are quantified in (log) fold change, we introduced soft interventions: “shift” $x\leftarrow f(\pi_{x})+c$, where $c\in\mathbb{R}$, and “scale” $x\leftarrow cf(\pi_{x})$, where $c$ is a positive scaling factor, with equal probability $c\lessgtr 1$. We also sampled interventions with multiple targets, to simulate off-target effects. Details are available in Section 4.2 and Appendix D.2.
Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995), and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable. We swept over the number of differential network layers (Figure 4) on synthetic data, and we used 3 layers for $h_{\text{cat}}$ and 2 layers for $h_{\text{diff}}$. Following Sea, we adopted hidden dimension $d=64$, the AdamW optimizer (Loshchilov & Hutter, 2019), learning rate 1e-4, batch size 16, and weight decay 1e-5. On the real data, where $N=1000$, we changed to a batch size of 1, decreased the learning rate to 5e-6, and finetuned the models with half precision (FP16). We trained both $h_{\text{cat}}$ and $h_{\text{diff}}$ models on synthetic data, but only finetuned the $h_{\text{diff}}$ architecture on the real data (memory constraint).
4 Experiments
@@
We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1). For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
@@ 4.1 Biological experiments
Datasets  We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al. (2022) and Nadig et al. (2024); as well as two Sci-Plex (Srivatsan et al., 2020) datasets (chemical perturbations) from McFaline-Figueroa et al. (2024). Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$. In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target (discussion in Appendix D.1).
Biologically, many perturbations do not affect the cell’s production of mRNA (they may affect other factors like protein activity, which transcriptomics does not measure). To ensure that $P_{\text{obs}}$ and $P_{\text{int}}$ are distinct, we filtered perturbations to those that induced over 10 differentially-expressed genes (statistically significant change, compared to control), of which the true target should be present. This excludes perturbations with insufficient cells (low statistical power), with minimal to no effect (uninteresting), and those that did not achieve the desired effect (low CRISPR efficiency). In the Sci-Plex data, several drugs have known off-targets, but we did not observe differential expression of any of them (adjusted p-value $\approx 1$).
@@
Evaluation  We consider two splits: seen and unseen cell lines. In the former, models may be trained on approximately half of the perturbations from each cell line, and are evaluated on the unseen perturbations. In the latter, we hold out one cell line at a time, and models may be trained on data from the remaining cell lines. To ensure that our train and test splits are sufficiently distinct, we cluster perturbations based on their log-fold change and assign each cluster to the same split (Figure 5). Since the Sci-Plex data are limited, we use them only as test sets without finetuning.
We limited the set of candidate targets to the top 1000 differentially expressed genes by log-fold change, per perturbation. If too few genes passed the significance threshold ($p<0.05$), we added additional candidates with the largest log-fold change until we reached a minimum of 100 genes. Genetic perturbations have highly specific effects compared to chemical perturbations, so we focused on the subset for which the intended target was not trivially identifiable as the gene with the largest log-fold change. [^1] Cdn predicts trivial perturbations nearly perfectly (Table 8).
@@
While perturbation target prediction appears to be a simple classification task, standard metrics are not well-suited for these real data. Not all genes are present in the baselines’ domain knowledge (e.g. isolated node in graph), so their effects as perturbations cannot be predicted. In addition, due to genetic redundancy, it is common for multiple perturbations to elicit similar responses (Kernfeld et al., 2023). An “incorrect” top 1 prediction may not necessarily reflect poor performance. Thus, we propose the following metrics.
@@
  • We compute the rank of the true target, normalized by the number of candidate genes. Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
  • We measure the similarity between the perturbation effect (log-fold change) of the top 1 predicted target and the true target with the Spearman rank correlation $\rho$ and the Pearson correlation $r$. These range from -1 to 1, where 1 is best.
@@
  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates. Recall ranges from 0 to 1, where 1 is best.
@@
Baselines  We compare to both algorithms that infer the effects of perturbations on cells (used as scoring functions, Equation 3), and algorithms that directly predict targets. All baselines were run with their official implementations and/or latest releases. For details, please see Appendix B.2.
Gears (Roohani et al., 2023) and GenePt (Chen & Zou, 2023) predict the effects of unseen genetic perturbations. Gears is a graph neural network that regresses log-fold change upon perturbation, using the Gene Ontology knowledge graph as an undirected backbone. GenePt is a set of large language model-derived embeddings that have been shown to achieve state-of-the-art performance (Märtens et al., 2024) with simple downstream models like logistic regression (used here). Both are trained per cell line.
PDGrapher (Gonzalez et al., 2024) is a graph neural network that predicts the perturbation targets of genetic or chemical perturbations, on seen or unseen cell lines, with the Human Reference Interactome (Luck et al., 2020) as the knowledge graph. To the best of our knowledge, PDGrapher is the only published deep learning baseline for the target prediction task on transcriptomic data. Therefore, we compare against additional statistical and naive baselines.
Dge stands for differential gene expression (Love et al., 2014), i.e. the process of identifying statistically significant changes in genes between settings. Specifically, we run the Wilcoxon ranked-sum test with Benjamini-Hochberg correction (Wilcoxon, 1945; Benjamini & Hochberg, 2000) between control and perturbed cells, as implemented by Wolf et al. (2018). Genes are ranked by adjusted p-value. Linear and Mlp take as input the mean expression of all perturbation targets, plus the top 2000 highly-variable genes (Wolf et al., 2018). They are trained to predict a binary label for each gene on each cell line independently.
@@
Table 1: Results on 5 Perturb-seq datasets. Top: Train on all cell lines jointly. Bottom: Leave one cell line out of training. Test sets (unseen perturbations) are identical in both settings. Cdn (synthetic) and Dge do not require training on real Perturb-seq data. Metrics, from left to right: normalized rank of ground truth, top 1 Spearman and Pearson correlations. Uncertainty quantification in Table 7. Runtimes in Table 11.
K562 (gw) K562 (es) RPE1 HepG2 Jurkat
Setting Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
Seen cell line Linear $0.51$ $0.18$ $0.23$ $0.48$ $0.22$ $0.25$ $0.49$ $0.42$ $0.52$ $0.51$ $0.35$ $0.43$ $0.50$ $0.20$ $0.23$
Mlp $0.45$ $0.18$ $0.21$ $0.49$ $0.08$ $0.11$ $0.53$ $0.32$ $0.41$ $0.48$ $0.33$ $0.40$ $0.45$ $0.24$ $0.29$
GenePt $0.52$ $0.29$ $0.32$ $0.42$ $0.16$ $0.19$ $0.41$ $0.37$ $0.46$ $0.31$ $0.32$ $0.40$ $0.51$ $0.31$ $0.36$
Gears $0.56$ $0.19$ $0.22$ $0.45$ $0.18$ $0.20$ $0.49$ $0.35$ $0.44$ $0.54$ $0.32$ $0.40$ $0.50$ $0.22$ $0.26$
PdG $0.54$ $0.25$ $0.29$ $0.49$ $0.20$ $0.23$ $0.48$ $0.41$ $0.52$ $0.54$ $0.37$ $0.43$ $0.52$ $0.31$ $0.35$
Cdn (Perturb-seq) 0.92 0.72 0.74 0.95 0.81 0.82 0.92 0.67 0.76 0.94 0.69 0.75 0.98 0.83 0.84
Unseen cell line Dge $0.82$ $0.50$ $0.54$ $0.89$ $0.60$ $0.62$ $0.79$ $0.56$ $0.64$ $0.86$ $0.51$ $0.51$ $0.84$ $0.45$ $0.50$
PdG $0.39$ $0.10$ $0.13$ $0.44$ $0.24$ $0.28$ $0.54$ $0.38$ $0.48$ $0.48$ $0.34$ $0.41$ $0.57$ $0.31$ $0.35$
Cdn (synthetic) $0.79$ $0.46$ $0.49$ $0.80$ $0.59$ $0.62$ $0.82$ $0.61$ $0.71$ $0.84$ $0.56$ $0.63$ $0.87$ $0.64$ $0.67$
Cdn (Perturb-seq) 0.91 0.69 0.71 0.95 0.81 0.82 0.92 0.65 0.74 0.93 0.67 0.72 0.97 0.84 0.85
Figure 3: Recall at $p$ (normalized over number of candidate genes on 5 Perturb-seq datasets. For Cdn, (P) and (S) denote “Perturb-seq” and “synthetic” training data.
@@
Table 2: Rank of drug targets on chemical perturbation datasets (McFaline-Figueroa et al., 2024). A172 and T98G are unseen cell lines. Highlighted: Within top 100. Drug names and details in Appendix D.1.
@@
Dge $0.12$ $0.64$ $0.28$ $0.94$ $0.24$ $0.83$
PdG $0.72$ $0.35$ $0.39$ $0.06$ $0.34$ $0.42$
Cdn $0.14$ $0.81$ $0.01$ $0.89$ $0.93$ $0.53$
Cdn (no $\mu$) $0.88$ $0.44$ $0.58$ $0.06$ $0.54$ $0.65$
Results  Cdn consistently outperforms all baselines on the five Perturb-seq datasets, in both the seen and unseen cell line settings (Table 1). There is minimal drop in between seen and unseen cell lines, suggesting that the functional characteristics of genetic perturbations are highly transferrable, even to different data distributions. Notably, no other deep learning method exceeds simple statistical tests (Dge). This is more evident in Figure 3, in which Cdn achieves higher recall at $p$ at nearly all points. Finally, if we ablate the finetuning step (“synthetic” vs. “Perturb-seq”), we find that finetuning leads to significant improvements.
Chemical perturbations are much less specific than genetic perturbations (Figures 6 and 7), and it is less clear whether drug effects should be reflected in mRNA levels (via feedback mechanisms), as drugs act upon their protein products. Still, on the two chemical perturbation datasets, Cdn finds the true target within the top 100 candidates for 3/6 cases (Table 2). Due to the low performance of Dge on several cases, we hypothesized that gene-level changes may not be as predictive. Indeed, an ablation of Cdn, trained without node-level statistics, improves where Dge fails (“no $\mu$”).